OpenStudy (2yoututu11):

Between what two consecutive integers does lie? A. 4 and 5 B. 5 and 6 C. 6 and 7 D. 7 and 8

OpenStudy (anonymous):

Does what lie?

OpenStudy (2yoututu11):

\[\sqrt{35}\]

OpenStudy (anonymous):

B: 5 & 6

OpenStudy (anonymous):

sqrt(36) is 6, so sqrt(35) must be smaller than 6. sqrt(25) is 5, so sqrt(35) is bigger than five :-)

OpenStudy (2yoututu11):

thank you all

OpenStudy (anonymous):

B. Because 5 squared is 25 and 6 squared is 36, it must lie somewhere between those two.

OpenStudy (anonymous):

or this Between what two consecutive integers does Square Root Of 35 lie? --------------------------------------… Let sqrt(35) = x, then x^2 = 35 But; 25 < x^2 = 35 < 36 So; sqrt(25) < sqrt(35) < sqrt(36) or 5 < sqrt(35) < 6 --------------------------------------… Hope this helps!